OpenStudy (anonymous):

Which expression is equivalent to this one? –7(5 + 10) A. (–7) · 5 + (–7) · 10 B. (–7) · 5 + 10 C. (–7) + 5 + 10 D. (–7) · 5 · 10

jigglypuff314 (jigglypuff314):

distributive property: (a) (b + c) = (a)·b + (a)·c so (-7) (5 + 10) =

OpenStudy (anonymous):

Which expression is equivalent to this one? 6(x – 17) A. 6x – 17 B. 6x · (–17) C. 6x – 6 · 17 D. 6 + x – 17

jigglypuff314 (jigglypuff314):

same idea as before :) distributive property: a(b - c) = ab - a·c so 6(x–17) =

OpenStudy (anonymous):

6x17?

OpenStudy (anonymous):

or no

OpenStudy (anonymous):

subtracting x -17 right

jigglypuff314 (jigglypuff314):

yes there is a subtraction :) so the end result would be something like 6x - 6 times 17

OpenStudy (anonymous):

Rewrite the expression by using grouping symbols. y • 4 + y • z A. y (4 + z) B. z(y • 4) C. y • (4 + y) • z D. 4(y + z)

jigglypuff314 (jigglypuff314):

same idea and now just reversed :) so distributive property: a·b + a·c = a(b + c) like take the thing that they have in common and put it out front :)

OpenStudy (anonymous):

so its c

jigglypuff314 (jigglypuff314):

not quite :) for example if it was y·7 + y·x then it would become y (7 + x)

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

i think i got answer

OpenStudy (anonymous):

lol never mind

jigglypuff314 (jigglypuff314):

lol well you can take a guess ;)

OpenStudy (anonymous):

a :o

jigglypuff314 (jigglypuff314):

Correct! :D :D :D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Evaluate the expression without using a calculator. 63 · 8 A. 484 B. 494 C. 504 D. 638

jigglypuff314 (jigglypuff314):

what is 3 times 8 = ?

OpenStudy (anonymous):

21

jigglypuff314 (jigglypuff314):

try again? :)

OpenStudy (anonymous):

hold up

OpenStudy (anonymous):

keep bored messing up i got it

OpenStudy (anonymous):

i know the answer just keybored messingup

OpenStudy (anonymous):

24

jigglypuff314 (jigglypuff314):

ok :) then 6 times 8 = ?

OpenStudy (anonymous):

48

jigglypuff314 (jigglypuff314):

then 48 + 2 = ? (the "2" is from the "24")

OpenStudy (anonymous):

50

jigglypuff314 (jigglypuff314):

so you get 504 the drawing tool isn't working for me right now... but this would be the basic concept http://www.wikihow.com/Do-Long-Multiplication

OpenStudy (anonymous):

oh cool

OpenStudy (anonymous):

last question

OpenStudy (anonymous):

What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

jigglypuff314 (jigglypuff314):

remember a few problems before when I showed you for example if it was y·7 + y·x then it would become y (7 + x) then what would 5•x + 5•7 become? :)

OpenStudy (anonymous):

mhmm

OpenStudy (ajanijones):

Your answer's C.

OpenStudy (anonymous):

to What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8 ?

OpenStudy (arabpride):

Distributive Property...

jigglypuff314 (jigglypuff314):

5(8 + 7) = 5 • x + 5 • 7 distribute 40 + 35 = 5x + 35 subtract 35 from both sides -35 -35 40 = 5x divide both sides by 5 so x = 40 divided by 5 = (simplify)

OpenStudy (anonymous):

um guys

OpenStudy (anonymous):

i got a 100%

jigglypuff314 (jigglypuff314):

you're welcome :P

OpenStudy (anonymous):

thanks Jigglypuff314

OpenStudy (anonymous):

ok guys

OpenStudy (anonymous):

i have one more test will you guys help me

OpenStudy (arabpride):

Good JOb @jigglypuff314 !!